### v1

# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora. However, training LLMs on human-written text entails significant risk of privacy and copyright violations, which demands an efficient machine unlearning framework to remove knowledge of sensitive data without retraining the model from scratch. While Gradient Ascent (GA) is widely used for unlearning by reducing the likelihood of generating unwanted information, the unboundedness of increasing the cross-entropy loss causes not only unstable optimization, but also catastrophic forgetting of knowledge that needs to be retained. We also discover its joint application under low-rank adaptation results in significantly suboptimal computational cost vs. generative performance trade-offs. In light of this limitation, we propose two novel techniques for robust and cost-efficient unlearning on LLMs. We first design an Inverted Hinge loss that suppresses unwanted tokens by increasing the probability of the next most likely token, thereby retaining fluency and structure in language generation. We also propose to initialize low-rank adapter weights based on Fisher-weighted low-rank approximation, which induces faster unlearning and better knowledge retention by allowing model updates to be focused on parameters that are important in generating textual data we wish to remove. Through extensive experiments on the Training Data Extraction Challenge dataset using the GPT-Neo family of LLMs, we find that our approach can efficiently remove unwanted information with negligible deterioration in reasoning as well as generative capabilities.
1 Introduction
Recently, Large Language Models (LLMs) have demonstrated substantial performance gains in downstream tasks as their model size and the amount of training data increase (Zhao et al., 2023). This has prompted many research groups to prioritize the collection of more data for pretraining them and the development of more larger models than before (Brown et al., 2020; Chowdhery et al., 2023; Rae et al., 2021; Smith et al., 2022). However, this approach has introduced significant privacy concerns, as these models tend to memorize data indiscriminately Carlini et al. (2021; 2022). For example, Personally Identifiable Information (e.g., names, phone numbers, and email addresses) can be easily extracted from the LLMs (Carlini et al., 2021). Additionally, OpenAI is currently facing several lawsuits related to alleged copyright infringement within their pretraining text corpus (Grynbaum & Mac, 2023). In response to these challenges and in light of the recent GDPR legislation on the right to be forgotten (Voigt & Von dem Bussche, 2017; Rosen, 2011; Villaronga et al., 2018), machine unlearning for LLMs has emerged as a critical and rapidly growing research field (Yao et al., 2023; Si et al., 2023a).
Exact unlearning can be achieved by filtering out sensitive information and retraining from scratch. However, this process is highly resource-intensive and inefficient, especially when data deletion requests occur serially rather than in large batches. As shown in Figure 1, this necessitates the development of approximate unlearning frameworks, where the goal is to remove knowledge of specific data instances from a pretrained model. Consequently, several methods for knowledge unlearning have been proposed (Wang et al., 2023; Chen & Yang, 2023; Jang et al., 2023). Among them, Jang et al. (2023) proposed a simple approach that fine-tunes LLMs using Gradient Ascent (GA) on the forgetting data, while also introducing several metrics to assess its effectiveness. They demonstrate that GA not only can be used for successful knowledge unlearning but also minimizes catastrophic forgetting of other knowledge within the LLMs. Nevertheless, GA still has significant drawbacks, such as the issue of training not converging due to unbounded loss and high unlearning cost caused by full fine-tuning for the entire LLM model.
Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it lose knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion. After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other data points in the retain set ${\mathcal{D}}_{r}$ as well as preserve its previously acquired reasoning and generative capabilities.
On the other hand, LoRA (Hu et al., 2021) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning for downstream tasks. The core idea behind LoRA is to freeze the pretrained LLM’s weights and train rank decomposition matrices within each layer during adaptation to downstream tasks. This approach not only significantly reduces the number of trainable parameters but also achieves performance comparable to full fine-tuning. Recently, Biderman et al. (2024) experimentally demonstrates that LoRA acts as a powerful regularization mechanism that prevents it from reaching the same performance levels as full fine-tuning. As a result, it results in less forgetting in retained knowledge. From this perspective, many of LoRA’s benefits are particularly advantageous in practical unlearning scenarios: 1) it significantly reduces the tuning cost for unlearning, and 2) it effectively prevents forgetting, making it a valuable method for achieving the important goal of retaining knowledge. Nonetheless, its application to LLM unlearning remains unexplored until now.
In this paper, we propose a novel approach for robust and efficient unlearning for LLMs by introducing a novel loss function and leveraging LoRA with a new initialization. First, we analyze the derivatives of the existing GA and highlight its shortcomings: 1) dispersed gradients causing unnecessary forgetting, and 2) an unbounded loss that fails to converge. To address these issues, we propose the Inverted Hinge loss (IHL). From the derivative analysis, we show that the IHL resolves all the problems of GA, enabling faster and robust unlearning. Second, we explore the intersection between LLM unlearning and cost-efficient fine-tuning using LoRA. We find that LoRA suffers from significantly suboptimal trade-offs between its downstream performance after unlearning and its ability to generate information requested for deletion. To achieve robust and efficient unlearning with LoRA, we introduce a new LoRA Initialization via Weighted Low-Rank Approximation. Extensive experiments on the LM-Eval and Dialogue benchmarks using Neo-GPT, Phi, and Llama2 LLMs show that our proposed Inverted Hinge Loss combined with LoRA achieves robust and cost-efficient unlearning than GA (See Figure 2).
Figure 2: Compute cost for successful unlearning vs. post-unlearning downstream performances of GPT-Neo-125M. Each point represents a different forget set and LoRA rank (if used). Left: Accuracy averaged across 9 classification tasks (higher is better). Middle: F1 score averaged across 4 dialogue generation tasks (higher is better). Right: Perplexity on the validation set of the Pile dataset (lower is better). Dashed lines indicate the performances of the model prior to unlearning. Unlearning with vanilla LoRA leads to significant loss in performance compared to full-parameter unlearning. However, our proposed method using both the Inverted Hinge Loss and FLoRA performs competitively to full-unlearning in all three aspects while enjoying high parameter- and cost-efficiency of LoRA.
2 Related Work
Machine Unlearning. The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pre-trained model to discard information acquired from a specific subset of data. Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2021; Mehta et al., 2022; Chundawat et al., 2022; Cha et al., 2024). Recently, its significance has grown notably in Large Language Models (LLMs) due to its pivotal role in managing unintended memorization intrinsic to LLMs (Si et al., 2023b; Yao et al., 2024).
First, Wang et al. (2023) introduces the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets. Chen & Yang (2023) proposes the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model frozen. While these approaches demonstrates impressive unlearning results across various datasets, they are hindered by significant drawbacks, such as the necessity to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) and the need to finetune models using them (Wang et al., 2023). Parameter merging-based methods recently garner attention as another category of unlearning for LLMs (Wu et al., 2023; Ilharco et al., 2022). Experimental results demonstrate that simple task arithmetic can facilitate cost-efficient unlearning for LLMs. However, these methods also exhibit limitations, such as achieving a relatively weak level of unlearning (Ilharco et al., 2022) or incurring higher computational costs to detect privacy neurons (Wu et al., 2023).
In contrast, following an unlearning approach similar to that used in image classification (Golatkar et al., 2020; Cha et al., 2024), Jang et al. (2023) leverages Gradient Ascent (GA) to maximize the next-token prediction loss for a sequence of tokens from the forget data. Unlike in image classification, they demonstrate that LLMs do not suffer from forgetting of retained knowledge when using GA on the forget data. Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks.
Parameter-Efficient Fine-Tuning. Fine-tuning pretrained LLMs to adapt towards specific downstream tasks and instructions comes at an intractable computational cost due to the large model size and complexity. To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023). Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2021; Zhang et al., 2023; YEH et al., 2023; Kopiczko et al., 2023; Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter. A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM. While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposes to initialize LoRA weights using the principal singular vectors and values of the linear weights. Inspired by this work, we propose to extract Fisher information from the data requested for unlearning and perform weighted low-rank approximation to prepare a LoRA initialization that induces faster unlearning and better preservation of general knowledge
3 Proposed Method
In this section, we introduce the notations and provide preliminaries on unlearning through Gradient Ascent (GA) and low-rank adaptation (LoRA). We then analyze the gradient properties of GA and highlight its inherent issues. Building on this analysis, we propose the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA. Finally, we present the idea of employing LoRA with a novel initialization strategy, thereby enabling cost-efficient unlearning.
3.1 Preliminaries
Problem and notation. Given a sequence of $T$ tokens ${\bm{x}}=(x_{1},x_{2},\dots,x_{T})$, a language model (LM) models the likelihood of the sequence via next token prediction: $p_{\theta}({\bm{x}})=\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{<t})$. After pretraining, we assume that an end-user has requested to delete a subset of the training set $\mathcal{D}_{f}\subset\mathcal{D}$, which we refer to as the forget set. Also, we consider a retain set $\mathcal{D}_{r}$ that represents the auxilary dataset which contains general knowledge the model should not forget after unlearning (e.g., Wikitext dataset (Merity et al., 2016)).
Gradient Ascent. Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, which leads to a simple yet effective baseline of gradient ascent (GA) (Cha et al., 2024; Jang et al., 2023). Unlike the usual gradient descent, GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
$\mathcal{L}_{GA}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$
Note that in practice, the log-likelihood is implemented in terms of cross-entropy losses on incremental tokens. Thus, GA tunes the model towards maximizing the cross-entropy, the value of which is unbounded above by nature, making optimization ill-posed and unstable.
Low-Rank Adaptation. Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank, LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter. In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes
$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}$ (1)
During fine-tuning, the original weight ${\bm{W}}$ is kept frozen and only the low-rank factors ${\bm{A}}$ and ${\bm{B}}$ are updated via gradient descent. To ensure that the initial attachment of LoRA adapters does not alter the output of the LLM, LoRA defaults to initializing ${\bm{A}}$ with a Kaiming-uniform distribution (He et al., 2015) and ${\bm{B}}$ as the zero matrix. Note that after fine-tuning, LoRA adapters can simply be merged with the original weights ${\bm{W}}^{\prime}={\bm{W}}+{\bm{B}}{\bm{A}}$, thereby avoiding any additional latency during inference. In the context of unlearning, this feature of LoRA naturally promotes a layer of security as it prevents adversarial agents from tracking previous records of unlearning just by inspecting the model architecture.
3.2 Preliminary Results
Despite its wide use in domain adaptation and instruction tuning, LoRA is not yet explored under the task of LLM unlearning to the best of our knowledge. Therefore, we first share empirical results from low-rank adapting LLMs using gradient difference (Maini et al., 2024; Liu et al., 2022a) as our objective to motivate our approach. For this experiment, we follow previous work (Jang et al., 2023) unlearn 32 text sequences from the Training Data Extraction Challenge from GPT-Neo-1.3B model pretrained on the Pile dataset (Gao et al., 2020). More information on the details on the setup can be found in Section 4 and the Appendix.
Figure 2 shows the results. Notably, vanilla LoRA suffers from lack of plasticity and ends up failing to sufficiently unlearn ${\mathcal{D}}_{f}$ within 20 epochs. When running more unlearning epochs or increasing the learning rate for sufficient unlearning, the model loses its previously acquired reasoning and generative capabilities, as shown in the significant decrease in LM-eval and Dialogue performances. In the remainder of this section, we present two techniques towards making LLM unlearning viable while enjoying the compute-efficiency of LoRA.
3.3 Inverted Hinge Loss: Novel Loss Function for Unlearning in LLMs
Motivation: Inherent Limitation of GA. We analyze the inherent issues of GA from the perspective of its gradients. The output layer of a language model is a softmax layer that outputs probabilities over the vocabulary. Let ${y}_{t}$ be the logits (pre-softmax activations) produced by the LLM model for the $t$-th token, and let $V$ be the vocabulary size. The probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$ where $y_{t}^{(x_{t})}$is the logit corresponding to the true token $x_{t}$ and $y_{t}^{(v)}$ is the logit corresponding to the $v$-th token in the vocabulary. When we use $\mathcal{L}_{GA}$ for unlearning for LLMs, the gradient of the log-probability with respect to the logits is:
$\frac{\partial\log(p_{\theta}(x_{t}|x_{<t}))}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
Given this, the derivative of $\mathcal{L}_{GA}({\bm{x}})$ with respect to the logits at each $t$ becomes $\nabla_{y_{t}^{(v)}}\mathcal{L}_{GA}({\bm{x}})={\partial\log(p_{\theta}(x_{t}|x_{<t}))}/{\partial y_{t}^{(v)}}$. From this derivative of GA, we can confirm its unlearning mechanism and inherent issues. First, when $v=x_{t}$, GA decreases the prediction score corresponding to the true token while increasing the prediction score for all other tokens. This demonstrates that GA performs unlearning by encouraging the model to predict tokens different from the true token.
However, despite the successful experimental results of GA (Jang et al., 2023) in LLMs, our gradient analysis reveals its inherent issues. When $v=x_{t}$, it increases the prediction scores of all tokens except the true token. As a result, GA has problems: 1) the loss not converging and becoming unbounded, and 2) unnecessary additional forgetting due to the increase in logits for all other tokens, 3) a large vocabulary size induces gradients to spread out across all other logits, making gradient updates for unlearning inefficient.
Inverted Hinge Loss (IHL). Considering the limitations of GA discussed above, our motivation is to develop a simple yet novel loss function for unlearning in LLMs. The new loss function aims to effectively achieve unlearning by decreasing the prediction score of the true token while increasing the prediction scores of only a minimal number of other tokens. Inspired by the hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge Loss (IHL) as shown below:
$\mathcal{L}_{\text{IHL}}({\bm{x}})=\max\left(0,1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))\right)$
Considering the probability $p_{\theta}(x_{t}|x_{<t})$ given by the softmax function, the derivative of $\mathcal{L}_{\text{IH}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}}.\end{cases}$
Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IH}}({\bm{x}})\neq 0$. Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IH}}({\bm{x}})=0$. The detailed process for deriving this is provided in the Appendix A
The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly demonstrates how the Inverted Hinge loss addresses the shortcomings of GA in knowledge unlearning for LLMs. First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with the opposite sign). This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly and intensively but also prevents spreading out of gradients for unlearning. During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$. Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease. This not only prevents unnecessary forgetting but also results in a bounded loss form.
3.4 FLoRA: Novel LoRA Initialization for LLM Unlearning
Motivation. Aside from reduing the number of trainable parameters and thus reducing the memory requirements of LLM fine-tuning, LoRA serves as a potent form of regularization. While it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios (Biderman et al., 2024). Under the task of LLM unlearning, however, our objective strictly requires to completely remove knowledge of ${\mathcal{D}}_{f}$, and the strong stability from LoRA imposes a strong burden: vanilla LoRA unlearning requires large number of iterations through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in our Figure 2). Inspired by previous work on PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are important to ${\mathcal{D}}_{f}$ but not so important to ${\mathcal{D}}_{r}$ would benefit the unlearning process in terms of both convergence and post-unlearning performance. We hence design a novel LoRA initialization technique that measures the Fisher information on each weight ${\bm{W}}$ targeted for LoRA, and perform weighted low-rank decomposition from which we use the principal components to initialize adapter weights ${\bm{A}}$ and ${\bm{B}}$.
Parameter Importances via Fisher Information. The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information dataset ${\mathcal{D}}$ provides on model parameters $\theta$. More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq. 3.4). However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead which, in the context of LLMs, can be computed as the average of squared gradients propagated from the language modeling objective (i.e. cross entropy loss):
${\bm{F}}_{\theta}=\mathbb{E}_{\mathcal{D}}\left[\left(\dfrac{\partial}{\partial\theta}\log p{\color[rgb]{0,0,0}{}_{\theta}}({\mathcal{D}}|\theta)\right)^{2}\right]\approx\dfrac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}\in{\mathcal{D}}}\left(\dfrac{\partial}{\partial\theta}\mathcal{L}_{\text{LM}}({\bm{x}};\theta)\right)^{2}\eqcolon\hat{{\bm{F}}}_{\theta},$ (2)
where $\mathcal{L}_{\text{LM}}$ is the next-token prediction loss of LM, such as $\mathcal{L}_{\text{LM}}({\bm{x}};\theta)=\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$. Within our LLM unlearning setup, a high empirical Fisher information measured with ${\mathcal{D}}_{f}$ indicates that the next-token prediction loss on ${\mathcal{D}}_{f}$ leads to large absolute gradients on the parameter under concern, and we consider such parameters to be important in generating sequences in ${\mathcal{D}}_{f}$.
Let $\hat{{\bm{F}}}_{\bm{W}}^{f}$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp. $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$). Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{rel}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to first identify parameters that are important specifically for ${\mathcal{D}}_{f}$, but unimportant to ${\mathcal{D}}_{r}$. While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus choose to use the relative Fisher information matrix rather than $\hat{{\bm{F}}}^{f}$.
Fisher-weighted LoRA Initialization (FLoRA). Intuitively, initializing LoRA weights to be parameters that are important in generating ${\mathcal{D}}_{f}$ would benefit unlearning as gradients would only modify parameters important to ${\mathcal{D}}_{f}$ while keeping all other parameters intact. Given the relative importance $\hat{{\bm{F}}}_{\bm{W}}^{rel}$ for each target parameter ${\bm{W}}$, we propose to initialize LoRA adapter weights as the solution to the following problem of Weighted Low-Rank Approximation (WLRA):
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\sum_{i,j}\left([\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}({\bm{W}}-{\bm{B}}{\bm{A}})_{i,j}\right)^{2}$ (3)
Note that when all weights equal one, this problem reduces down to a standard low-rank approximation problem, of which the solution can easily be computed via rank-$r$ SVD. In general, however, this minimization problem does not have a closed-form solution and thus requires iterative optimization (Srebro & Jaakkola, 2003). While we may resort to iterative methods towards initializing LoRA weights in a fine-grained manner, this would undermine the gains in efficiency due to employing LoRA. Thus, we tackle a simplified version of WLRA instead by assuming that parameters in each row of ${\bm{W}}$ share the same importance, for which we use the square-root of the row-wise sum of the relative Fisher information matrix as
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\left\|\text{diag}\left((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\ \bm{\cdot}\ \mathds{1})^{\frac{1}{2}}\right)({\bm{W}}-{\bm{B}}{\bm{A}})\right\|_{2}$ (4)
where $\mathds{1}\in\mathbb{R}^{k}$ is the all-one vector. $\text{diag}(\cdot)$ and $\bm{\cdot}$ denote the vector diagonalization function and the matrix-vector product, respectively. Unlike general WLRA, the row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
Given the solution, we use the solution ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights. To ensure that the output is not perturbed after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}{\bm{A}}$. In summary, our Fisher-Weighted initialization extracts parameters that are important only to ${\mathcal{D}}_{f}$ and not to ${\mathcal{D}}_{r}$ such that LoRA tuning can be focused towards erasing knowledge of ${\mathcal{D}}_{f}$.
3.5 Final Loss Function for Unlearning
Following initialization with FLoRA, unlearning is performed on the model $\Theta=\theta\cup\theta_{FLoRA}$, which consists of the original weights $\theta$ and the LoRA weights for each layer $\theta_{FLoRA}=\{{\bm{A}}^{*}_{l},{\bm{B}}^{*}_{l}\}_{l=1}^{L}$, where $L$ represents the number of layers where FLoRA is applied. During unlearning process, $\theta$ remains frozen, and only updates to the FLoRA weights $\theta_{FLoRA}$ are made. Furthermore, when performing unlearning on the forget dataset $D_{f}$, we incorporate the auxiliary retain set $D_{r}$, which contains general knowledge (e.g., the Wikitext dataset), as a default setting, following the gradient difference approach (Liu et al., 2022a; Maini et al., 2024). The final loss function using the proposed Inverted Hinge Loss is defined as follows:
$\operatorname*{minimize}_{\theta_{FLoRA}}\sum_{{\bm{x}}_{r}\in D_{f},{\bm{x}}_{f}\in D_{r}}\mathcal{L}_{\text{IH}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (5)
In practice, training (unlearning) for the LLM model involves sampling corpora from each dataset, $D_{f}$ and $D_{r}$, and minimizing the final loss function through stochastic gradient descent.
4 Experiments
In this section, we first perform unlearning experiments based on the Training Data Extraction Challenge (§4.1), followed by ablation and analytical results (§4.2). All experiments were conducted on a remote GCP server equipped with 16 NVIDIA A100 Tensor Core GPUs.
Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset. Note that, in all experiments using LoRA, GD is used as the default loss function. The "+IHL" results refer to experiments where GA within GD is replaced with the proposed IHL (i.e., using Eqn.(5)), and "+FLoRA" indicates the results where FLoRA is used instead of LoRA.
Model Method Params. Epochs EL10 MA Reasoning Dialogue Pile
(%) (%$\downarrow$ (%$\downarrow$ (Acc)$\uparrow$ (F1$\uparrow$ (PPL)$\downarrow$
Before - - 30.9 77.4 43.4 9.4 17.8
GA 100.0 17.2 1.0 27.4 39.9 2.6 577.8
GPT-Neo GD 4.6 0.7 24.9 42.4 5.9 54.2
125M LoRA 1.6 8.6 0.3 20.6 40.8 2.5 129.4
+ IHL 11.4 0.4 22.7 41.9 6.0 32.9
+ FLoRA 6.0 0.3 23.9 42.2 10.1 24.0
Before - - 67.6 92.2 49.8 11.5 11.5
GA 100.0 13.8 1.9 30.4 49.7 8.5 15.8
GPT-Neo GD 12.8 2.2 30.9 48.4 12.7 10.8
1.3B LoRA 0.8 19.3 1.7 31.4 45.0 9.7 31.8
+ IHL 20.0 1.7 44.6 47.1 10.2 14.9
+ FLoRA 13.0 0.5 29.6 48.3 12.1 14.7
Before - - 70.4 93.4 52.3 11.5 10.4
GA 100.0 10.8 1.6 31.0 51.9 11.1 17.9
GPT-Neo GD 8.0 0.7 28.3 44.0 12.7 17.9
2.7B LoRA 0.7 14.0 0.1 20.4 45.9 6.7 61.1
+ IHL 17.8 0.0 26.7 49.6 8.5 22.2
+ FLoRA 10.3 0.1 28.5 49.6 10.7 16.0
4.1 Training Data Extraction Challenge
Experimental Setup.
Assuming a scenario where an adversary attacks a pretrained LLM to extract text samples previously used in training, the Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) that are found to be easily extractable. For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$. For the retain set ${\mathcal{D}}_{r}$, we use the validation subset of WikiText as it contains factual world knowledge that we wish to maintain within the LLM.
For TDEC experiments, we consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using three different forget sets. For this experiment, we use a fixed learning rate of 5e-5 when full finetuning, and 2e-4 when using LoRA and FLoRA. For all model sizes, we use a LoRA (and FLoRA) rank of $r=16$ as a default. We set the default to applying LoRA to the attention matrices for Query (Q) and Value (V), as well as to the Feedforward Network (FFN).
Following previous work (Jang et al., 2023), we measure the unlearning efficacy via two metrics. The ${\bm{n}}$-gram Extraction Likelihood (ELn) measures the $n$-gram overlap between the ground truth sequence in ${\mathcal{D}}_{f}$ and the output generated by the model. The Memorization Accuracy (MA) measures the token-wise accuracy of the LLM on ${\mathcal{D}}_{f}$. More details on these metrics are introduced in Section B of the Appendix. After each unlearning epoch, we measure EL10/AM and consider the model to have successfully unlearned ${\mathcal{D}}_{f}$ if both values are smaller than those measured using a held-out validation set. Once unlearning is finished, we evaluate the unlearned model on various downstream benchmarks to ensure the LLM maintains its previously acquired language modeling capabilities after unlearning. To assess its reasoning capabilities, we average accuracies across 9 different classification datasets LAMBADA (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), COPA (Gordon et al., 2012), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), PiQA (Bisk et al., 2020), MathQA (Amini et al., 2019), PubmedQA (Jin et al., 2019). To measure generative performance, we also measure the F1 score averaged over four dialogue generation datasets, Blended Skill Talk (Smith et al., 2020), Empathetic Dialogues Rashkin et al. (2019), Wizard of Internet (Komeili et al., 2021), Wizard of Wikipedia (Dinan et al., 2018). Lastly, we measure the perplexity on the validation subset of the Pile (Gao et al., 2020).
Baseline.
For experiments on knowledge unlearning in LLMs, we select using two baseline algorithms, Gradient Ascent (GA) (Jang et al., 2023) and Gradient Difference (GD) (Liu et al., 2022a; Maini et al., 2024), which rely solely on the original language model while utilizing a forget dataset $D_{f}$ only or both a forget dataset $D_{f}$ and an auxiliary dataset $D_{r}$, respectively. For all experiments, we use the same forgetting criteria for ELn and MA as in  Jang et al. (2023) and report the results after performing unlearning until these criteria are met.
Figure 3: Results from unlearning 32 examples from Training Data Extraction Challenge. Each row represents the average performance in each set of LM capabilies: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better). Each column shows results under different model size: 125M, 1.3B, and 2.7B. Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed. X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics. Solid lines indicate the performance of different methods averaged only across successful unlearning trials. The dashed lines show respective performances prior to unlearning. Unlearning with vanilla GD not only leads to significant loss in performance, but also fails to unlearn sufficiently even with large LoRA ranks. Replacing the negative cross-entropy loss with our hinge loss (IHL) shows significant gains in reasoning and generation capabilities, but still fails to unlearn a number of example sets. Using NCE but with Fisher-weighted LoRA initialization (FLoRA) notably increases the rate of successful unlearning, but at the cost of significant loss in overall LM capability. Using both IHL and FLoRA best minimizes post-unlearning performance degradation in all three capabilities under successful unlearning.
Results.
Table 1 shows the evaluation results of full-parameter unlearning and unlearning with LoRa and FLoRA. From our experimental results, we can identify the following findings. First, for all model sizes, using GD for unlearning not only achieves the forgetting criteria in fewer epochs compared to using GA alone, but it also better preserves previously acquired knowledge. Specifically, for the 125M and 1.3B models, knowledge unlearning with GA leads to a significant deterioration in PPL on the Pile dataset, whereas GD not only overcomes this issue but also achieves improved Dialogue (F1) performance compared to Before (e.g., GPT-Neo 1.3B). Second, we observe that simply applying GD loss with LoRA fails to achieve successful unlearning across all model sizes. Despite the fact that only about 0.7% to 1.6% of the total parameters are updated for unlearning when using LoRA, its simple application actually causes forgetting of previously acquired knowledge. Finally, when using the proposed IHL along with FLoRA, it achieves the forgetting criteria in fewer epochs while better preserving previously acquired knowledge. Although using IHL with LoRA alone maintains previously acquired knowledge better than the baseline, it requires relatively more epochs to meet the forgetting criteria. However, when combined with FLoRA, not only is the number of epochs required to meet the forgetting criteria significantly reduced, but overall, previously acquired knowledge is preserved more successfully.
Additionally, we conduct further experiments with different ranks for LoRA and FLoRA, with results and detailed explanations provided in Figure 3 and its caption. From all experiments so far, we could confirm that simply applying LoRA to GD does not lead to successful unlearning. However, using the proposed IHL in combination with FLoRA demonstrates better cost-efficiency than the existing GA and GD methods, while also better preserving previously acquired knowledge during unlearning.
Figure 4: Results from unlearning 32 examples from Training Data Extraction Challenge using LoRA for GPT-Neo 2.7B. Note that each color denotes Unlearning with vanilla GD, Replacing the negative cross-entropy loss with our hinge loss (IHL), Using NCE but with Fisher-weighted LoRA initialization (FLoRA), and Using both IHL and FLoRA. Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed. X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics.
4.2 Analysis
What modules do we need to adapt?
Figure 4 presents experiments where LoRA or FLoRA is applied to various weight matrices, including those for Query (Q), Value (V), Key (K), Output (O) in the self-attention module, and the FeedForward Network (FFN). While the original LoRA paper indicates that applying LoRA to Q and V yields superior performance on downstream tasks (Hu et al., 2021), our experiments show that applying LoRA to Q and V alone is insufficient to meet the unlearning criteria in all cases. Notably, applying LoRA to the FFN results in much more successful unlearning in cases of using FLoRA. Furthermore, applying FLoRA to QVFFN in conjunction with IHL not only achieves superior unlearning results but also better preserves previously acquired knowledge. Based on these findings, we adopt the application of LoRA and FLoRA to QVFFN as the default setting.
Cost-efficiency of the proposed method.
To compare unlearning costs, we evaluate the FLOPs for each method and the results are presented in Figure 2. While using vanilla LoRA offers significant advantages in terms of unlearning cost (i.e., FLOPs), as confirmed by our previous experiments, it results in substantial performance losses compared to full-parameter unlearning. In contrast, combining the proposed Inverted Hinge Loss with FLoRA not only achieves the best performance but also leverages the cost advantages of LoRA, demonstrating that this approach enables cost-efficient unlearning.
5 Concluding Remarks
In this paper, we address limitations of existing unlearning techniques (i.e., Gradient Ascent) by introducing the novel Inverted Hinge Loss (IHL) and leveraging LoRA with a novel initialization (i.e., Fisher-weighted LoRA Initialization (FLoRA)). Our findings reveal that IHL overcomes key drawbacks associated with Gradient Ascent (GA), such as dispersed gradients and unbounded loss, leading to faster and more robust unlearning. Furthermore, our approach of combining FLoRA with IHL not only reduces the cost of unlearning but also enhances the preservation of previously acquired knowledge. Through experiments, we show that our method achieves superior performance in knowledge unlearning compared to the baseline method.
Appendix A Derivative Analysis for the Inverted Hinge Loss Function
The function $p_{\theta}(x_{t}|x_{<t})$ represents a probability distribution that indicates the likelihood of $x_{t}$ taking a specific token $x_{t}$ given the previous tokes $x_{<t}$. This probability is expressed using the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$, where $y_{t}^{(v)}$ denotes the score for the $v$-th token in the vocabulary. To differentiate this function with respect to $y_{t}^{(x_{t})}$, we rewrite $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{Z}$ where $Z=\sum_{v=1}^{V}\exp(y_{t}^{(v)})$ is the normalization constant.
We differentiate this function with respect to $y_{t}^{(k)}$ considering two cases: 1) $k=x_{t}$ and 2) $k\neq x_{t}$. For the first case, we can get the following by using the chain rule:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(x_{t})}}=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(\frac{\exp(y_{t}^{(x_{t})})}{Z}\right)=\frac{1}{Z}\frac{\partial\exp(y_{t}^{(x_{t})})}{\partial y_{t}^{(x_{t})}}-\frac{\exp(y_{t}^{(x_{t})})}{Z^{2}}\frac{\partial Z}{\partial y_{t}^{(x_{t})}}$
Here, $\frac{\partial\exp(y_{t}^{(x_{t})})}{\partial y_{t}^{(x_{t})}}=\exp(y_{t}^{(x_{t})})$ and $\frac{\partial Z}{\partial y_{t}^{(x_{t})}}=\exp(y_{t}^{(x_{t})})$. Therefore, it becomes:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(x_{t})}}=\frac{\exp(y_{t}^{(x_{t})})}{Z}-\frac{\exp(y_{t}^{(x_{t})})^{2}}{Z^{2}}=p_{\theta}(x_{t}|x_{<t})-p_{\theta}(x_{t}|x_{<t})^{2}=p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))$
For the first case, using the chain rule again, we get:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(k)}}=\frac{\partial}{\partial y_{t}^{(k)}}\left(\frac{\exp(y_{t}^{(x_{t})})}{Z}\right)=-\frac{\exp(y_{t}^{(x_{t})})}{Z^{2}}\frac{\partial Z}{\partial y_{t}^{(k)}}$
where $\frac{\partial Z}{\partial y_{t}^{(k)}}=\exp(y_{t}^{(k)})$. Therefore,
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(k)}}=-\frac{\exp(y_{t}^{(x_{t})})\exp(y_{t}^{(k)})}{Z^{2}}=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(k|x_{<t})$
Thus, we can summarize them as below:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))&\text{if }v=x_{t}\\ -p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
Based on the derivative of $p_{\theta}(x_{t}|x_{<t})$ above , we can calculate the derivative of $\mathcal{L}_{\text{IH}}$. Firstly, for convenience, we define $p_{t}=p_{\theta}(x_{t}|x_{<t})$ and $\hat{p}_{t}=\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$. The loss function can be rewritten as:
$\mathcal{L}_{\text{IH}}({\bm{x}})=\max\left(0,1+p_{t}-\hat{p}_{t}\right)$
To calculate the derivative of $\mathcal{L}_{\text{IH}}$, we need to consider three cases: 1) when $v=x_{t}$, 2) when $v=v^{\star}$ where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$, 3) when $v\neq x_{t}$ and $v\neq v^{\star}$. In the case where $01+p_{t}-\hat{p}_{t}>$ 0, using the derivative of $p_{\theta}(x_{t}|x_{<t})$ mentioned earlier, the derivative of $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ is as follows:
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(x_{t})}}$ $=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
$=p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))+p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v^{\star}|x_{<t})$
$=p_{\theta}(x_{t}|x_{<t})\left(1-p_{\theta}(x_{t}|x_{<t})+p_{\theta}(v^{\star}|x_{<t})\right)$
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v^{\star})}}$ $=\frac{\partial}{\partial y_{t}^{(v^{\star})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
$=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})(1-p_{\theta}(v^{\star}|x_{<t}))$
$=-p_{\theta}(v^{\star}|x_{<t})\left(1-p_{\theta}(v^{\star}|x_{<t})+p_{\theta}(x_{t}|x_{<t})\right)$
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v)}}$ $=\frac{\partial}{\partial y_{t}^{(v)}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
$=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v|x_{<t})+p_{\theta}(v^{\star}|x_{<t})\cdot p_{\theta}(v|x_{<t})$
$=p_{\theta}(v|x_{<t})\left(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})\right)$
Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $1+p_{t}-\hat{p}_{t}\leq 0$. In summary, the derivatives of the loss function $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ for the three cases are:
$\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$
Appendix B Evaluation Metrics
How to measure success of unlearning? Following previous work Jang et al. (2023); Tirumala et al. (2022), we empirically measure the success of unlearning using two metrics, Extraction Likelihood (EL) and Memorization Accuracy (MA), which we briefly discuss below.
After unlearning each sequence ${\bm{x}}=(x_{1},\dots,x_{T})\in\mathcal{D}_{f}$, the Extraction Likelihood (EL) is measured as the $n$-gram overlap between the ground truth sequence ${\bm{x}}$ and the output of the model after unlearning.
$\textsc{Overlap}_{n}({\bm{a}},{\bm{b}})=\dfrac{\sum_{{\bm{c}}\in\textsc{$n$-gram}({\bm{a}})}\mathds{1}\{{\bm{c}}\in\textsc{$n$-gram}({\bm{b}})\}}{|\textsc{$n$-gram}({\bm{a}})|}$ (6)
$\textsc{EL}_{n}({\bm{x}})=\dfrac{\sum_{t=1}^{T-n}\textsc{Overlap}_{n}\left(f_{\theta}(x_{<t}),x_{\geq t}\right)}{T-n}$ (7)
The Memorization Accuracy (MA) measures the token-wise memorization of the LM $p_{\theta}$.
$\textsc{MA}({\bm{x}})=\dfrac{\sum_{t=1}^{T}\mathds{1}\{\operatorname*{arg\,max}_{x}p_{\theta}(x|x_{<t})=x_{t}\}}{T-1}$ (8)
Given these two metrics, we flag successful unlearning when the average EL and MA on $\mathcal{D}_{f}$ goes below the EL and MA values measured on the validation set unseen during training. In our experiments we measure EL with 10-grams, which results in the following early stopping criterion.
$\dfrac{1}{\mathcal{D}_{f}}\sum_{{\bm{x}}\in\mathcal{D}_{f}}\textsc{EL}_{10}({\bm{x}})\leq\dfrac{1}{\mathcal{D}_{\text{val}}}\sum_{{\bm{x}}\in\mathcal{D}_{\text{val}}}\textsc{EL}_{10}({\bm{x}})\;\;\;\;\text{and}\;\;\;\;\dfrac{1}{\mathcal{D}_{f}}\sum_{{\bm{x}}\in\mathcal{D}_{f}}\textsc{MA}({\bm{x}})\leq\dfrac{1}{\mathcal{D}_{\text{val}}}\sum_{{\bm{x}}\in\mathcal{D}_{\text{val}}}\textsc{MA}({\bm{x}})$